OpenStudy (anonymous):

Simplify: (2^2 + 3^4)^-2

OpenStudy (callisto):

2^2 = 2x2 = ... 3^4 = 3x3x3x3 =... (2^2 + 3^4)^-2 = \(\frac{1}{(2^2 + 3^4)^2}\) =...

OpenStudy (luigi0210):

\[x^-1=\frac{ 1 }{ x }\]